Suppose that an engineer determines that she can use the formula -t^2+14 less than or equal to P to estimate when the price of carbon fibre will be P dollars per kilogram or less in t years from the present. when will the carbon fibre be available at $10/kg or less?
Answer by ikleyn(52775)

.
Suppose that an engineer determines that she can use the formula -t^2+14 less than or equal to P to estimate when the price
of carbon fibre will be P dollars per kilogram or less in t years from the present. when will the carbon fibre be available at $10/kg or less?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

To determine "t", you need to solve the equation

-t^2 + 14 = 10.

----->

-t^2 = 10 - 14,

-t^2 = -4,

t^2 = 4,

t = 2.

Answer.  When ?  - In two years from now.